if a spring has a constant k and the spring is cut in three parts what will be the constant for each of the part of the spring
1
Expert's answer
2013-02-01T07:35:25-0500
F = k*Δl
F- force
Δl - deformation
k = F/Δl
Δl ~ l
Therefore,
ki = k*l/li
l - length of full spring
li - length of parts (l1+l2+l3 = l)
For example, if li = l/3 -> ki=3*k &